# No Vulnerability Data, No Problem: Towards Predicting Mean Time To Remediate In Open Source Software Dependencies

### v2

@@ 5.3. Metrics Implementation
In the database, we compute if the resolved dependency version matches the highest available version of the dependency during each interval and mark each interval accordingly, as shown in the “updated” column in Table. 2. Similarly, we then compute if the resolved dependency version for each interval is vulnerable to any security advisory even though a fixed version is available and mark the interval accordingly in the “remediated” column in Table. 2.
5.4. Statistical and Correlation Tests
To answer RQ1, we claim that if MTTUdep and MTTRdep are statistically equivalent, then MTTUdep can serve as a proxy for MTTRdep. Since we did not expect our samples to follow a normal distribution, we initially selected the non-parametric Mann-Whitney test. However, we also applied the t-test to ensure a more conservative analysis, as our data is not ordinal. Both tests were conducted using the Python library scipy, and the Wilcoxon test was also implemented using the same package. To assess normality, we used the Kolmogorov-Smirnov test.
Our initial hypotheses focused on the equivalence of data distributions. For these, following the approach of Papotti et al. (papotti_acceptance_2024, ), we employed the TOST (Two One-Sided Test) approach, originally proposed by Schuirmann et al. (schuirmann_hypothesis-testing_1981, ) and commonly used in pharmacological and food sciences to evaluate equivalence within a specified range (guidance_statistical_2001, ; meyners_equivalence_2012, ). Other research also used the TOST test in their analysis (shelton_rosita_2021, ; labunets2018no, ; labunets2017graphical, ; palheiros2024does, ). According to TOST, two distributions, $x$ and $y$, are considered equivalent if $x\cdot\delta<y<x\cdot 1/\delta$, where $\delta=0.8$. The TOST method can use either Mann-Whitney test or t-test as its underlying direction test. We reported the results of the t-test because it was more conservative than the Mann-Whitney test.
Additionally, to measure the association between MTTUdep and MTTRdep, we calculated both the Pearson correlation coefficient and Spearman’s rank correlation coefficient (similar to Cox et al. (cox_measuring_2015, )). High ($>0.7$) or moderate coefficient values ($>0.5$) would suggest a strong to moderate positive relationship (ratner2009correlation, ). Such findings would also support the notion that MTTUdep could effectively serve as a proxy for MTTRdep.
@@ 5.5. Package Characteristics Analysis
For analyzing eight package-level characteristics in RQ1, we use correlation tests to estimate if there is any correlation between these characteristics and packages’ corresponding MTTUdep and MTTRdep values. Specifically, we test a hypothesis about the association between the following independent variables and their continuous dependent variables (MTTUdep and MTTRdep):
Contributors Count ($H_{1}$). We hypothesize packages with fewer contributors are less likely to have updated dependencies since they have less capacity for maintenance and dependency management.
@@
Dependents Count ($H_{2}$). We hypothesize packages with fewer dependents are less likely to have updated dependencies. Fewer dependents may lead to fewer feature updates, fewer bug fixes, and fewer version releases, which in turn may result in less updated dependencies.
Dependency Count ($H_{3}$). We hypothesize packages with fewer dependencies are more likely to have updated dependencies. We expect fewer dependencies to be more manageable, and thus, these projects may have more updated dependencies.
Version Count ($H_{4}$). We hypothesize packages with fewer number of version releases are less likely to have updated dependencies. Since these packages have less available versions, they may pay less attention to their dependency management.
Major Version Count ($H_{5}$). We hypothesize packages with fewer major version releases are more likely to have updated dependencies. Since these packages have fewer major versions to maintain, they may pay more attention to their dependency management.
@@
Package Age ($H_{6}$). We hypothesize packages with lower ages are more likely to have updated dependencies. Lower package age may mean the development team is more proactive in dependency management since the package is not old or mature enough. Lower package age may also mean that the dependencies do not have scope for publishing many newer versions, which in turn may mean more updated dependencies in these packages.
@@
SourceRank ($H_{7}$). The Package SourceRank score indicates the package quality, popularity, and community metrics calculated in libraries.io dataset (saini_investigating_2020, ; sun_using_2023, ). This metric depends on several factors, such as the presence of a README file, license, following SEMVER, recent updates, and the number of contributors. We hypothesize packages with lower SourceRank scores are less likely to have updated dependencies. Since these packages are of low quality, they may not have a lot of dependents or contributors, which in turn may result in less updated dependencies.
@@
Forks And Stars Count ($H_{8}$). We hypothesize packages with fewer forks and stars are less likely to have updated dependencies. Fewer forks and stars may mean fewer people are using and looking into these packages, which in turn may mean less activity, less version releases, and less updated dependencies in these packages.
@@ 5.6. Regression Analysis
Correlation results indicate whether two independent variables are associated with each other or not. However, correlation does not establish causation. To identify such a cause-and-effect relationship using mathematical modeling, we opt for applying a technique of multilinear regression analysis. The multilinear regression model takes independent variables (e.g., package characteristics) and one dependent variable (MTTUdep or MTTRdep) and gives results on the relationship between each independent variable and the dependent variable. This analysis provides valuable insights on predicting MTTUdep or MTTRdep using the package’s characteristics. This analysis also results in $p$-values and coefficients, indicating which package characteristics might have a significant impact on MTTUdep or MTTRdep.
@@
This process will produce a $p$-value for each independent variable to indicate whether the relationship between this variable and the outcome is statistically significant. Finally, we consider the variables with $p$-value $<0.05$ as significant variables, and the package characteristics associated with these variables are identified as key package characteristics. We use statsmodels package of Python to conduct the correlation tests and to build the regression model.
6. Results
6.1. RQ1: Can MTTUdep serve as a proxy for MTTRdep?
We statistically assess the relationship between MTTUdep and MTTRdep to determine if MTTUdep can effectively serve as a proxy for MTTRdep in  RQ1.
We used TOST (Two One-Sided Test) to investigate the equivalence of MTTUdep and MTTRdep. The single one-sided tests are done with the Mann-Whitney test. The results of the TOSTs are reported in Table 3. We also visualized the TOST for MTTUdep and MTTRdep from all three ecosystems in Fig. 2.
Table 3. TOSTs Results
Ecosystem MTTUdep ×$\times 0.8$ $<$ MTTRdep $<$ MTTUdep ÷$\div 0.8$
[…]
The green area corresponds to $x\cdot\delta<y<x\cdot 1/\delta$ where $\delta=0.8$, $x$ is MTTUdep, and $y$ is MTTRdep.
Figure 2. Distribution behind TOST MTTUdep vs. MTTRdep.
The results of the TOSTs show statistical equivalence between MTTUdep and MTTRdep ($p<0.05$) within the given range. Further, we also ran correlation analyses between MTTUdep and MTTRdep, in which both Spearman and Pearson correlation coefficients are $0.5$. The correlation analysis supports the results of our TOST tests as we can conclude that MTTUdep and MTTRdep are moderately and positively correlated since all correlation coefficients are $>=0.5$.
This result confirms that as the MTTUdep increases for a package, it is highly probable that the MTTRdep also increases (and vice versa), supporting the use of MTTUdep as a proxy for MTTRdep when we do not have enough vulnerability data to compute MTTRdep.
Takeaway 1: MTTUdep can be used as a proxy for MTTRdep since MTTUdep and MTTRdep are statistically equivalent (and moderately and positively correlated) in all three observed ecosystems.
6.2. RQ1: How do other package characteristics influence MTTUdep and MTTRdep?
6.2.1. Correlation Analysis.
We tested eight package-level characteristics that might influence MTTUdep and MTTRdep. We present the correlation matrix with the packages’ characteristics and MTTUdep and MTTRdep data in Figure 3. The correlation heatmap provides insight into the relationships between various factors and their potential impact on MTTUdep and MTTRdep. The results indicate weak correlations between MTTUdep and the other factors.
Contributors Count ($H_{1}$): We hypothesized that packages with fewer contributors would be less likely to maintain updated dependencies. From the heatmap, we can see that the contributors count show minimal correlations with MTTUdep (-0.04) and MTTRdep (-0.07), suggesting a limited influence on dependency management.
Dependents Count ($H_{2}$): We hypothesized that packages with fewer dependents are less likely to have updated dependencies. The results indicate a negligible correlation between dependents count and MTTUdep or MTTRdep. This suggests that the number of dependents does not play a substantial role in determining the timeliness of dependency updates.
Dependency Count ($H_{3}$): We hypothesized packages with fewer dependencies to have more updated dependencies. The results show almost no correlation between dependency count and MTTUdep or MTTRdep. This implies that the number of dependencies in a package is not associated with dependency management efficiency.
Version Count ($H_{4}$): We hypothesized that packages with fewer version releases are less likely to have updated dependencies. However, the correlation matrix revealed no meaningful relationship between the number of version releases and either MTTUdep or MTTRdep. This suggests that the frequency of version releases does not significantly impact how quickly dependencies are updated.
Major Version Count ($H_{5}$): We hypothesized that packages with fewer major version releases are more likely to have updated dependencies. However, the correlation matrix revealed no meaningful relationship between the number of major version releases and either MTTUdep or MTTRdep. This suggests that the number of available major versions to maintain does not significantly impact how quickly dependencies are updated.
Package Age ($H_{6}$): We hypothesized that packages with lower ages are more likely to have updated dependencies. However, the correlation matrix invalidated this hypothesis with a weak correlation score ($0.32$). This suggests that as packages become mature with age, they tend to keep dependencies less updated. However, this conclusion should be treated with caution since our metrics might be age-sensitive (more discussion on this in Section 7).
SourceRank ($H_{7})$. We hypothesized that low quality packages (e.g., packages with lower SourceRank scores) are less likely to have updated dependencies. The correlation matrix supports this hypothesis for MTTRdep with a weak correlation score ($-0.22$), which means low-quality packages have higher MTTRdep values. However, low quality packages do not necessarily have higher MTTUdep values ($-0.06$ coefficient).
Forks And Stars Count ($H_{8}$): We hypothesized that packages with fewer forks and stars are less likely to have updated dependencies. The results reveal weak to moderate positive correlations between stars (0.65) and forks (0.72) with maintainers count, indicating that more popular packages may have more active development. However, their relationships with MTTUdep and MTTRdep remain insignificant, suggesting that forks and stars do not directly translate into faster dependency updates.
Overall, our results indicate weak or negligible correlations between the tested factors and dependency management efficiency as measured by MTTUdep and MTTRdep. These findings suggest that the analyzed factors may not strongly influence how updated dependencies are, and further exploration of other variables or non-linear relationships is recommended.
Takeaway 2: Package characteristics (except package age and SourceRank score) have negligible correlations with dependency updatedness. Older packages are associated with higher MTTUdep, and popular and better quality packages (higher SourceRank score) are associated with lower MTTRdep.
6.2.2. Regression Analysis (MTTUdep As Dependent Variable).
Interpreting the multilinear regression model’s characteristics, we found 0.073 as R-squared value when MTTUdep was the dependent variable. R-squared value 0.073 indicates that 7.3% variation in the dependent variable can be explained by the model. Although the model is statistically significant (p-value ¡ 0.05 from the F-test), lower R-squared value indicates that other factors, beyond the nine dependent variables, also substantially impact MTTUdep. The F-statistic of this model is large (1399), and Prob(F-statistic) is 0.00. This indicates that at least one of the independent variables (or predictors) has a non-zero relationship with MTTUdep.
We then look into the coefficients and p-values associated with each of the independent variables. The coefficient indicates the expected change in the dependent variable for a one-unit change in one independent variable while holding other independent variables constant. The p-value associated with the t-statistic indicates whether the independent variable is statistically significant in explaining the variation in the dependent variable. A lower p-value (¡ 0.05) would indicate statistical significance.
We have found that some independent variables have positive coefficients, which indicates an increase in one of these variables would result in an increase in our dependent variable, MTTUdep. Independent variables with positive coefficients are contributors count (0.0204), stars (0.0035), dependencies count (2.3027), and package age (0.1257). The dependencies count has the largest positive coefficient, which means that packages with more dependencies need longer times to update their dependencies. In addition, older packages tend to have longer MTTUdep as indicated by the 0.1257 coefficient, justifying our hypothesis ($H_{3}$) on dependencies count.
We also found some independent variables have negative coefficients. Negative coefficients indicate that an increase in that independent variable would result in a decrease in the dependent variable, MTTUdep. Independent variables with negative coefficients are dependents count (-0.0013), forks (-0.0085), SourceRank (-24.3203), and major versions count (-0.9068). SourceRank, which indicates package quality and popularity, shows strongly negative coefficients. Higher SourceRank indicates an overall better project. The model predicts if SourceRank for a package increases by 1 (moving to a “better” rank), MTTUdep drops by 24 days. Moreover, packages that have released more major versions tend to have a lower MTTUdep. Finally, all dependent variables except dependent count were found to be statistically significant (p-value ¡ 0.05).
Takeaway 3: Packages with a lot of dependencies or older packages are more likely to have higher MTTUdep. Also, popular and better quality packages (with higher SourceRank scores) are more likely to have lower MTTUdep.
6.2.3. Regression Analysis (MTTRdep As Dependent Variable).
For MTTRdep as the dependent variable, the multilinear regression model results in 0.097 as R-squared value. R-squared = 0.097 indicates that the model explains 9.7% of the variation in MTTRdep. A lower R-squared value indicates that other unmodeled factors play an important role in determining MTTRdep, which is similar to our observations in MTTUdep. The F-statistic of this model is 246.8, and Prob(F-statistic) is 0.00. This indicates the model is statistically significant and at least one of the dependent variables (or predictors) has a non-zero relationship with MTTRdep.
We then look into the coefficients and p-values associated with each of the independent variables. We found that forks, stars, SourceRank, dependencies count, major version count and package age are statistically significant (p-value ¡ 0.05). On the other hand, the contributors count, and dependents count are not statistically significant (p-value ¿ 0.05) in modeling MTTRdep.
As a predictor, forks (coefficient 0.0070) and stars (coefficient 0.0014) indicate a positive but small effect. More forks and stars are associated with slightly higher MTTRdep. Dependencies count (coefficient 0.5028), and package age (coefficient 0.1239) show a moderate positive effect on MTTRdep. Additionally, the major version count (coefficient -2.4954) shows a moderate negative effect, whereas SourceRank (coefficient -34.8183) shows a strong negative effect on MTTRdep. This observation is similar to our previous observation with MTTUdep. According to the model’s prediction, an increase in SourceRank by 1 (i.e., a higher SourceRank indicates a better package) would result in a 34.8-day reduction in MTTRdep.
Takeaway 4: Similar to MTTUdep, older packages are more likely to have higher MTTRdep. Also, popular and better quality packages (with higher SourceRank scores) are more likely to have lower MTTRdep. Packages with higher major versions are more likely to have lower MTTRdep.
6.3. RQ1: How long do packages have outdated dependencies, and how long do packages have vulnerable dependencies?
@@ 6.2. RQ1: How do other package characteristics influence MTTUdep and MTTRdep?
In RQ1, we empirically analyze the MTTUdep and MTTRdep metrics for the three ecosystems. We choose violin plot instead of box-plot since violin plot shows everything box-plot shows, e.g., medians, ranges, variability, and violin plot’s shape shows the density of the data similar to a density estimation plot (hintzeViolinPlotsBox1998, ).
@@
Fig. 4 shows a violin plot of MTTUdep in npm, PyPI, and Cargo. All the plots are right-skewed, indicating that most packages have a low MTTUdep. For instance, $50\%$ npm package has MTTUdep of less than $174$ days. In terms of distribution width, Cargo has a compact distribution, and npm is the most widely spread (Cargo ¡ PyPI ¡ npm). Also, every plot has a long tail, indicating that every ecosystem has some packages that do not update their dependencies for a long time (max MTTUdep $=4296$ days).
Fig. 5 shows a violin plot visualizing the distribution across MTTRdep in days for Cargo, npm, and PyPI packages. The first, second (mean), and third quantiles are shown as , , and , respectively. The overall pattern is similar to MTTUdep. The relative distribution width and the right-skewed nature of MTTRdep is similar to MTTUdep. Cargo and PyPI have concentrated distributions, with most MTTRdep (e.g., interquartile range) being less than $219$ and $281$ days. Conversely, npm shows a more spread distribution than Cargo and PyPI, indicating variability in MTTRdep in npm.
We could compute MTTUdep for $163,207$ ($117,129$ npm, $42,777$ PyPI, and $3,301$ Cargo) packages, and MTTRdep for $22,513$ ($17,263$ npm, $5,158$ PyPI, and $92$ Cargo) packages. This corroborates our initial motivation for conducting this study, the lack of vulnerability data.
Takeaway 5: Even without potentially abandoned packages, a subset of packages show extreme delays in keeping their dependencies and vulnerable dependencies updated.

### v3

@@ 5.3. Metrics Implementation
In the database, we compute if the resolved dependency version matches the highest available version of the dependency during each interval and mark each interval accordingly, as shown in the “updated” column in Table 2. Similarly, we then compute if the resolved dependency version for each interval is vulnerable to any security advisory, even though a fixed version is available, and mark the interval accordingly in the “remediated” column in Table 2.
5.4. Proxy Analysis
To answer RQRQ3:: , we apply some statistical tests on MTTUdep and MTTRdep data to analyze whether a variable (MTTUdep) can be used as a substitute/ proxy for another (MTTRdep) from the literature.
(1) TOST (Two One-Sided Test). Schuirmann et al. (schuirmann_hypothesis-testing_1981, ) proposed an equivalence test, known as TOST (Two One-Sided Test), in bioequivalence studies to determine if a treatment (e.g., a drug) can be used as a substitute for another. This method then was used in pharmacological/ food science (guidance_statistical_2001, ; meyners_equivalence_2012, ), medical research (mascha2011equivalence, ), and later also adopted to software engineering and security (papotti_acceptance_2024, ; shelton_rosita_2021, ; labunets2018no, ; labunets2017graphical, ; palheiros2024does, ). According to TOST, two distributions, $x$ and $y$, are considered equivalent if $x\cdot\delta<y<x\cdot 1/\delta$, where $\delta=0.8$. We report TOST with Mann-Whitney tests as the underlying difference tests.
(2) Regression with Wald Test. Several works (wooldridge2010econometric, ; mahnken2014evaluating, ) proposed the use of regression to identify proxy. A proxy should have (1) a statistically significant slope ($p_{value}$ ¡ 0.05); (2) a normal distribution for random error with a mean of zero and small variance. Additionally, Montgomery et al. (montgomery2000measuring, ) used (3) a Wald test on the proxy variable coefficient. Based on these, we ran an Ordinary Least Squares (OLS) regression using the statsmodel library in Python, with a Wald test.
(3) Sensitivity Analysis. Seltzer (seltzer2021perilous, ) proposed sensitivity analysis for proxy analysis, with $GFI$ (Goodness-of-Fit Index), $AGFI$ (Adjusted Goodness-of-Fit Index), and $NFI$ (Normal Fit Index). The indices should be ¿ 0.90 for an acceptable fit (marsh1996assessing, ). We ran the sensitivity analysis using the semopy library in Python.
(4) Correlation Analysis. High correlation is one indication of a proxy (frost1979proxy, ). We calculated both the Pearson correlation coefficient (as used by Oh et al. (oh2019identification, )) and Spearman’s rank correlation coefficient (similar to Cox et al. (cox_measuring_2015, )). High (>0.7>0.7) or moderate coefficient values (>0.5>0.5) would suggest a strong to moderate positive relationship (ratner2009correlation, ). We utilize the scipy.stats package from Python to calculate these correlations.
@@ 5.5. Package Characteristics Analysis
For analyzing nine package-level characteristics in RQRQ4:: , we use correlation tests to estimate if there is any correlation between these characteristics and packages’ corresponding MTTUdep and MTTRdep values. Specifically, we test a hypothesis about the association between the following independent variables and their continuous dependent variables (MTTUdep and MTTRdep):
@@
($H_{1}$) Contributors Count.. We hypothesize that packages with fewer contributors are less likely to have updated dependencies since they have less capacity for maintenance and dependency management. ($H_{2}$) Dependents Count. We hypothesize that packages with fewer dependents are less likely to have updated dependencies. Fewer dependents may lead to fewer feature updates, fewer bug fixes, and fewer version releases, which in turn may result in less updated dependencies. ($H_{3}$) Dependency Count. We hypothesize that packages with fewer dependencies are more likely to have updated dependencies. We expect fewer dependencies to be more manageable, and thus, these projects may have more updated dependencies.
@@
($H_{4}$) Version Count. We hypothesize that packages with fewer version releases are less likely to have updated dependencies. Since these packages have fewer available versions, they may pay less attention to their dependency management. ($H_{5}$) Major Version Count. We hypothesize that packages with fewer major version releases are more likely to have updated dependencies. Since these packages have fewer major versions to maintain, they may pay more attention to their dependency management. ($H_{6}$) Package Age. We hypothesize that packages with lower ages are more likely to have updated dependencies. Lower package age may mean the development team is more proactive in dependency management since the package is not old or mature enough. Lower package age may also mean that the dependencies do not have scope for publishing many newer versions, which in turn may mean more updated dependencies in these packages.
@@
($H_{7}$) SourceRank. The Package SourceRank score indicates the package quality, popularity, and community metrics calculated in libraries.io dataset (saini_investigating_2020, ; sun_using_2023, ). This metric depends on several factors, such as the presence of a README file, license, following SEMVER, recent updates, and the number of contributors. We hypothesize that packages with lower SourceRank scores are less likely to have updated dependencies. Since these packages are of low quality, they may not have a lot of dependents or contributors, which in turn may result in less updated dependencies.
@@
($H_{8}$) Forks Count. We hypothesize that packages with fewer forks are less likely to have updated dependencies. Fewer forks may mean fewer people are using and looking into these packages, which in turn may mean less activity, fewer version releases, and fewer updated dependencies in these packages.
($H_{9}$) Stars Count. We hypothesize that packages with fewer stars are less likely to have updated dependencies. Fewer stars may indicate fewer people are using and looking into these packages, which consequently may suggest less activity, fewer version releases, and fewer updated dependencies.
@@ 5.6. Regression Analysis
Correlation analysis measures pairwise relationships between two variables, but does not account for potential interactions between multiple variables. In contrast, a multilinear regression model allows evaluating the combined effect of all independent variables in predicting the dependent variable while controlling for the others. Moreover, strong relationships between two variables from correlation analysis might be influenced by the presence of other variables (confounding variables). A multilinear regression model controls for confounding effects, providing an understanding of the unique contribution of each independent variable in predicting the dependent variable. So, for RQRQ4:: , we use multilinear regression models that take independent variables (e.g., package characteristics) and one dependent variable (MTTUdep or MTTRdep) and give results on the relationship between each independent variable and the dependent variable. This analysis provides valuable insights into predicting MTTUdep or MTTRdep using the package’s characteristics. This analysis also results in $p_{value}$s and coefficients, indicating which package characteristics might have a significant impact on MTTUdep or MTTRdep.
@@
This process will produce a $p_{value}$ for each independent variable to indicate whether the relationship between this variable and the outcome is statistically significant. To control for family-wise type-I error inflation due to testing multiple dependent models (e.g., models that share the same dependent variable) together, we applied the Bonferroni correction (bonferroni1935, ) for the $p_{value}$ threshold to determine the statistical significance level. Since we test nine different characteristics, according to Bonferroni correction, the $p_{value}$ of a test needs to be <0.05/9or 0.0055<0.05/9\ or\ 0.0055 to be considered as significant ($p<0.0055$). Package characteristics with a significant test result are identified as key package characteristics. We use the statsmodels package of Python to conduct the correlation tests and to build the regression model.
6. Empirical Study Results
6.1. RQRQ2:: : How do packages in npm, PyPI, and Cargo differ in MTTU and MTTR?
@@
In RQRQ2:: , we empirically analyze the MTTUdep and MTTRdep metrics for the three ecosystems. We choose a violin plot instead of box plot since a violin plot shows everything a box plot shows, e.g., medians, ranges, variability, and the violin plot’s shape shows the density of the data similar to a density estimation plot (hintzeViolinPlotsBox1998, ). Fig. 3 shows a violin plot of MTTUdep in npm, PyPI, and Cargo. All the plots are right-skewed, indicating that most packages have a low MTTUdep. For instance, $50\%$ npm package has MTTUdep of less than $51$ days. Also, every plot has a long tail, indicating that every ecosystem has some packages that do not update their dependencies for a long time (max MTTUdep =2653=2653 days). In addition, interquartile ranges are small and comparable for the three ecosystems (Cargo with $1\sim 39$ days, npm with $4\sim 45$ days, and PyPI with $11\sim 54$ days).
Fig. 3 shows a violin plot visualizing the distribution across MTTRdep in days for Cargo, npm, and PyPI packages. The overall pattern is similar to MTTUdep. The relative distribution width and the right-skewed nature of MTTRdep are similar to MTTUdep. Interquartile ranges in MTTRdep are comparable for the three ecosystems (Cargo with $6\sim 24$ days, npm with $10\sim 42$ days, and PyPI with $12\sim 45$ days), similar to MTTUdep. A smaller interquartile range indicates MTTUdep and MTTRdep data is less spread and less variable. We could compute MTTUdep for $163,207$ ($117,129$ npm, $42,777$ PyPI, and $3,301$ Cargo) packages, and MTTRdep for $22,513$ ($17,263$ npm, $5,158$ PyPI, and $92$ Cargo) packages. This corroborates our initial motivation for conducting this study: the lack of vulnerability data.